Suppose a ball rebounds 2/3 the distance it falls. If it is dropped from a height of 20 m, how far does it travel before coming to rest?
Answer by stanbon(75887)

Suppose a ball rebounds 2/3 the distance it falls. If it is dropped from a height of 20 m, how far does it travel before coming to rest?
-----------
20+ 20(2/3) + 20(2/3)^2 + ... + 20(2/3)^(n+1)
--------------
Geometric Series with a = 20 and r = (2/3)
-----
Sum = a/(1-r) = 20/(1/3) = 60 m
---------
Cheers,
Stan H.